Grade 12Algebra

In a shop there are 5 icecreams available. A child buys six icecreams. IN how many ways he can do so..?

Nandini
9 Years agoGrade 12

1 Answer

Arun
9 Years ago
Let no. of ice-creams bought of each type as a,b,c,d & e.
now a+b+c+d+e=6
thus the required no. of ways= 6+5-1C5-1 = 10C4 = 210